# Superpixel Tokenization for Vision Transformers: Preserving Semantic Integrity in Visual Tokens

### v1

@@ 4.2 Superpixel-aware Feature Aggregation
where $|\mathcal{C}_{k}|$ and $\mathbf{F}^{(h,w)}$ denote the number of pixels within superpixel $\mathcal{C}_{k}$ and the pre-aggregate feature at coordinate $(h,w)$, respectively.
Average pooling and max pooling extract complementary features from each superpixel. Average pooling captures the overall characteristics common in the superpixel, while max pooling identifies the most salient features that may be obscured by averaging alone. We concatenate $D/2$-dimensional vectors from each pooling method along the channel dimension, yielding a $D$-dimensional token $\mathbf{z}^{(k)}$:
@@ 5 Experiments
Size Model # Params. GMACs # tokens Top-1 Tiny DeiT 5.7M 1.26 196 72.2 SuiT 5.5M 0.79 100 72.2 1.26 171 75.3 1.44 196 75.7 Small DeiT 22.1M 4.61 196 79.8 SuiT 21.7M 3.64 132 79.8 4.60 172 80.5 5.20 196 80.9 Base DeiT 86.8M 17.6 196 81.8 SuiT 86.0M 16.4 160 81.8 17.6 173 82.0 19.7 196 82.1
Table 1: Comparison of classification performance on ImageNet-1K against DeiT. The performance of SuiT is reported with varying token counts. Underlined values indicate those that match DeiT’s GMACs or token counts.
@@ 5.1 Experimental Settings
We validate our tokenization method by applying it to ViTs, dubbed as SuiT. We experiment on three model scales, Tiny ($D=192$), Small ($D=384$), and Base ($D=768$). The values $d_{local}$ and $d_{pos}$ are both set to $D/4$.
We employ FastSLIC [4] as our superpixel algorithm due to its high speed. We use a stride of 2 in the convolutional layer of local feature extraction, which produces feature maps of 2×2$\times$2 × smaller size than the input. The positional feature map $\mathbf{F}_{pos}$ and superpixel index map $\mathcal{I}$ are scaled to the same size accordingly.
This design choice was based on the observation that using a stride of 1 offered no notable performance improvements while significantly increasing computational costs. Further details are described in Appendix.
5.2 Comparison to Baseline
As our approach focuses on tokenization without modifying the backbone network, we focus on comparison with our baseline, DeiT [50] on ImageNet-1K classification [12]. We mostly follow the experimental settings from DeiT, with minimal changes for fair comparison.
Table 1 compares the classification performance of SuiT to that of DeiT. For each model size, the table provides a detailed comparison across key aspects: classification accuracy, computational cost measured in GMACs, and the number of tokens processed. This adaptive inference capability arises from the flexibility to adjust the number of superpixel tokens, enabling users to customize the configuration based on specific needs. When achieving comparable performance, SuiT exhibits substantially higher efficiency than DeiT. At equivalent GMACs, SuiT delivers superior accuracy. Maintaining an identical number of tokens further boosts its accuracy,
Tiny Small Base
Quadformer [44] - 80.8 84.4
MSViT [16] 75.5 82.0 -
SuiT (Fine-tuned) 76.1 82.6 84.5
SPiT [1] - 75.0 80.4
SuiT (Scratch) 75.7 80.9 82.1
Table 2: Comparison with existing studies on tokenization. Experiment above reports the Top-1 accuracy on ImageNet-1K. The numbers for Quadformer-Tiny, MSViT-Base and SPiT-tiny are intentionally left blank, as they are unavailable.
@@ 5.2 Comparison to Baseline
DINO-ViT 72.6 78.6 93.0 97.1
DINO-SuiT 72.8 79.4 93.8 97.2
Table 3: Results of transfer learning different datasets in terms of top-1 accuracy results. SuiT outperforms ViT in most downstream tasks in both supervised and self-supervised settings. All results for DeiT and DINO are reproduced following the training recipes provided by the authors. : The reported Flowers102 performance could not be reproduced in the exact setting, so we compare with our reproduced results.
@@
5.3 Comparison to Existing Tokenization Methods
We evaluate our tokenization method against existing approaches under two scenarios used in existing tokenization studies. For a fair comparison, we follow the setup of  [44, 16], which fine-tunes a ViT pretrained on ImageNet-21K [43]. Additionally, we include SPiT  [1], a concurrent work leveraging superpixels for tokenization, which trains its model from scratch on ImageNet-1K. Both experiments are conducted on ImageNet-1K classification tasks.
As reported in Table 2, our method consistently outperforms existing tokenization methods across all scales of model size, showing its efficacy. On fine-tuning experiments following [44, 16], SuiT outperforms existing tokenization methods for fine-grained tokens, such as Quadformer [44] and MSViT [16].
In experiments for training from scratch following [1], SuiT consistently outperforms SPiT, another work on superpixel-based tokenization, by a noticeable margin. These results verify the excellence of our tokenization pipeline compared to prior work.
5.4 Transfer Learning
We verify the generalization ability of SuiT when transferred to various downstream tasks. We follow the standard fine-tuning protocol and default settings of DeiT [50] for fair comparison, by fine-tuning ImageNet-1K pre-trained SuiT to each downstream task specifically. We test our models on iNaturalist [22], Flowers102 [39], and StanfordCars [10]. Further details can be found in the Appendix.
Table 3 presents a comparison of the classification accuracy between DeiT and SuiT when fine-tuned on various downstream tasks. Across various tasks and model scales,  SuiT consistently achieves higher accuracy than DeiT. Notably, on iNaturalist2018, SuiT-Small outperforms DeiT-Small by 0.8%p, demonstrating its superior ability to effectively transfer learned features, particularly when ample data is available.
Figure 3: Self-attention map comparison between DINO-ViT and DINO-SuiT. All maps are visualized by averaging the [CLS] token’s self-attention from the final layer across all heads.
@@ 5.4 Transfer Learning
5.5 Self-supervised Learning
We further study the applicability of SuiT in self-supervised scenario. Specifically, we train SuiT on ImageNet-1K with DINO [9], a well-known self-supervised learning approach, chosen for its proven effectiveness and popularity.
5.5.1 Transfer Learning
After pre-training SuiT with DINO, we evaluate its performance when transferred to downstream tasks to confirm the quality of learned features. As shown in the lower section of Table 3, DINO-SuiT consistently surpasses DINO-ViT across all downstream tasks. This demonstrates that SuiT effectively supports self-supervised scenarios, further establishing its strength as a versatile pipeline for a wide range of vision tasks.
5.5.2 Zero-shot Salient Object Segmentation
It is well known that the attention map of ViTs trained with DINO effectively attends to object regions without explicit supervision [9]. As shown in Figure 3, while DINO-ViT’s self-attention maps are limited to square patch-level details, those of DINO-SuiT even capture finer structures of salient objects. To demonstrate that this fine-grained property synergistically enhances the salient object segmentation capabilities of DINO [9], we measure the segmentation performance on three benchmark datasets, ECSSD [57], DUTS [54] and DUT-OMRON [58] following the works of Wang et al. [56], Aasan et al. [1]. We conduct our experiments on the TokenCut framework [56], which is a normalized graph cut algorithm that extracts saliency mask from attention map.
Superpixel tokenization significantly improves DINO-ViT [9] performance across most metrics without post-processing, achieving results comparable to DINO-ViT with post-processing.
Notably, as shown in Table 4, our model exhibits minimal performance differences between post-processed and non-post-processed cases, unlike DINO-ViT. Compared to SPiT [1], SuiT demonstrates competitive performance across three datasets. However, SPiT [1] does not show its applicability to self-supervised methods such as DINO and is trained with larger resolution images. Qualitative comparisons in Figure 4 highlight that DINO-SuiT detects all salient objects, even in multi-object scenarios, whereas DINO-ViT shows limitations such as rectangular masks without post-processing and incomplete object detection.
@@ 5.5.2 Zero-shot Salient Object Segmentation
6 Analysis
6.1 Semantic Integrity of Superpixel Tokens
To verify the improved semantic preservation of our superpixel tokens, we perform K-means clustering [26] on tokens embedding space. We compare the clustering results of superpixel tokens with grid-based tokens.
As shown in Figure 5, our superpixel tokens are grouped based on semantic similarity rather than spatial proximity, unlike the patch tokens from DeiT, which lack consistency within clusters. For example, in the first row, tokens corresponding to the regions of ducklings are grouped together, whereas clusters from DeiT’s patch tokens fail to convey clear meaning. This analysis suggests that superpixel tokens are better at preserving semantics. Additional results are provided in the Appendix.
Figure 6: Class-identifiability analysis results of DeiT and SuiT. We show the proportion of class-identifiable [CLS] tokens and image tokens for SuiT and DeiT across all layers.
@@ 6.1 Semantic Integrity of Superpixel Tokens
Table 5: Experimental results of SuiT and DeiT on ImageNet-A and ImageNet-O. We report accuracy on ImageNet-A and AUPR on ImageNet-O, following  [21].
6.2 Class Identifiability Analysis
Vilas et al. [52] examined class-specific encoding in supervised ViTs by projecting intermediate representations through the trained classifier head into a class-specific embedding space, similar to the logit lens method in NLP [40, 6]. Building on this, we evaluate class-specific encoding across SuiT’s layers by projecting token representations through the classifier head and measuring the proportion of tokens correctly predicting the target class.
Figure 6 shows that SuiT’s classification token shows higher class identifiability in the earlier layers compared to DeiT. This can be attributed to the semantic preserving nature of superpixel tokenization, which aids in aggregating class-specific encoding in early layers. In the deeper layers, DeiT’s patch tokens exhibit class identifiability levels slightly lower than its classification token, with a marginal gap. This observation aligns with findings from previous studies [52, 53], which report that patch tokens in supervised ViTs tend to encode increasingly global information as layers deepen. In contrast, SuiT’s patch tokens maintain a lower proportion of class-identifiable features in the later layers, suggesting a reduced emphasis on global feature encoding. This divergence alludes that SuiT and DeiT adopt distinctive strategies to aggregate class-specific features at different network depths.
6.3 Enhanced Robustness
@@
In Table 5, SuiT consistently outperforms DeiT on both datasets across model variants.
These findings indicate that our superpixel tokenization improves model robustness in challenging scenarios by mitigating the model’s reliance on features that are easily learned but lack generalizability.
6.4 Ablation Studies
Local feats. PE Freqs. PE Injection Top-1
random init. concat+proj 75.7
random init. concat+proj 66.5
No PE 71.5
predefined concat+proj 73.0
random init. addition 75.2
random init. concat 73.8
Avg Std Softmax Max Top-1
75.7
73.2
75.5
Fail
75.5
Pre-aggregate Features
The following ablation studies on our proposed components are conducted using SuiT-Tiny. Table 6 highlights the importance of components in pre-aggregate feature extraction. First, using raw 3-channel RGB features leads to a significant performance drop, confirming the need for higher-dimensional local features. Second, removing positional information reduces performance but retains partial class prediction through local cues. Experiments also reveal that predefined positional frequencies under-perform compared to learned frequencies, emphasizing the need for vision-specific customization. Lastly, our results show that the combination of concatenation and projection more effectively integrate local and positional features than simple concatenation or addition used in prior studies [51, 14, 50].
Aggregation
Table 7 presents the analysis of our aggregation method, which employs cardinality-independent operations. Removing max pooling results in a notable performance drop, and removing average pooling causes a slight decrease. We also tested standard deviation as an alternative to max pooling, but it proved numerically unstable and unsuitable for training. Softmax pooling, which computes a weighted sum of values within each superpixel, showed promising results but slightly underperformed compared to max pooling. This may be due to the softmax operation suppressing prominent features crucial for recognition, leading to less discriminative representations.
7 Conclusion
In this work, we proposed a novel superpixel-based tokenization method for ViTs as an attempt to improve the visual tokenization process. By leveraging superpixels, the proposed tokenization obtains semantic-preserving tokens, overcoming the issue of mixed semantics in grid-based patches. To incorporate superpixels that are irregular in shape, size, and location, we proposed a two-stage pipeline consisting of pre-aggregate feature extraction and superpixel-aware feature aggregation. Extensive experiments demonstrated the effectiveness of our method on diverse tasks, e.g., supervised classification, transfer learning, and self-supervised learning. Additional analytical findings provided meaningful insights into the mechanism and emergent properties of superpixel tokenization. We hope our promising results inspire further exploration of token design in ViTs, fostering advancements in both model performance and interpretability.
Supplementary Material
A1 Implementation Details
A1.1 ImageNet-1K classification
A1.1.1 Training from scratch
The settings for experiments from scratch is presented in Table A2. We mostly follow the settings of DeiT [50] with minimal changes. The main difference to the DeiT setting is as follows: 1) to align with our main concept of using superpixels, we disable MixUp [61]. 2) In random erasing [64], the randomly erased region is filled with per-channel random color instead of per-pixel random color. 3) We use gradient clipping, which was known to be helpful in ViT [14], but later removed in DeiT [50]. Other than that, we slightly tweak hyper-parameters of learning rate, weight decay and stochastic depth. For models of all three scales, Tiny, Small and Base, we use a base learning rate of $48e-$ 4, which is scaled proportionally to the batch size, by ×$\times\frac{batchsize}{512}$. We use different weight decay and stochastic depth values for each scale. Most notably, our model uses a larger weight decay and drop rate for stochastic depth, especially in Base scale. This is because we found that our model tends to overfit quickly, compared to the baseline DeiT. We thus opt to use a stronger regularization term. Similar observations were made in Transfer Learning, which is also discussed in the Section  A1.2.
A1.1.2 Fine-tuning from Pre-trained weights
@@ 7 Conclusion
A1.2 Transfer Learning
@@
A1.3 Self-Supervised Learning
For DINO [9] pre-training, we follow the exact training hyper-parameters used to pre-train ViTs, without any tuning, extra training techniques or augmentations. We train DINO-SuiT for 300 epochs, and transfer the pre-trained model to downstream tasks following the settings mentioned in Section A1.2. Additional self-attention maps of DINO-SuiT can be found in Figure A1.
A1.4 Zero-shot Salient Object Segmentation
@@
A2 Semantic Integrity of Superpixel Tokens

### v3

@@ 4.2 Superpixel-aware Aggregation
where $|\mathcal{C}_{k}|$ and $\mathbf{F}^{(h,w)}$ denote the number of pixels within superpixel $\mathcal{C}_{k}$ and the pixel-level embeddings at coordinate $(h,w)$, respectively.
Average pooling and max pooling extract complementary embeddings from each superpixel. (5) Average pooling captures the overall characteristics common in the superpixel, while max pooling identifies the most salient features that may be obscured by averaging alone. We concatenate $D/2$-dimensional vectors from each pooling method along the channel dimension, yielding a $D$-dimensional token $\mathbf{z}^{(k)}$:
@@ 5 Experiments
Size Model Weight-Init. # Params. GMACs Top-1 (%)
Tiny DeiT [55] Random 5.7M 1.26 72.2
SPiT [1] - - -
SuiT 5.5M 1.44 75.7
MSViT [19] IN-21K - 0.98 75.5
Quadformer [49] - - -
SuiT 5.5M 1.44 76.1
Small DeiT [55] Random 22.1M 4.61 79.8
SPiT [1] - - 75.0
SuiT 21.7M 5.20 80.9
MSViT [19] IN-21K - 4.43 82.0
Quadformer [49] - 4.71 80.8
SuiT 21.7M 5.20 82.6
Base DeiT [55] Random 86.8M 17.6 81.8
SPiT [1] - - 80.4
SuiT 86.0M 19.7 82.3
MSViT [19] IN-21K - - -
Quadformer [49] - 17.7 84.4
SuiT 86.0M 19.7 84.5
Table 1: Comparison of classification performance on ImageNet-1K [12]. Following prior studies, we condcuted experiments on two different settings for fair comparisons. In the “Weight-Init.” column, “Random” refers to from-scratch training on ImageNet-1K [12], while “IN-21K” refers to training from ViT weights pre-trained on ImageNet-21K [48]. denotes that the method was initialized with IN-21K pre-trained + IN-1K fine-tuned ViT weights. Cells marked ‘-’ means that the value is not available.
Figure 3: Adaptive inference. Under the same image resolution (=224), DeiT [55] is forced to use a fixed # tokens (=196), whereas both SuiT and SPiT [1] can adaptively adjust the number of tokens during inference. Notably, across all token counts and model scales, SuiT consistently achieves the best performance compared to baselines under the same computational cost.
@@ 5.1 Experimental Settings
We validate our tokenization method by applying it to ViT, referring to the resulting model as SuiT. As baselines, we include DeiT [55] and three tokenization-focused approaches—SPiT[1], Quadformer[49], and MSViT[19]—while keeping the backbone unchanged.
We evaluate three model scales: Tiny $(D=192)$, Small $(D=384)$, and Base $(D=768)$, setting the local and positional feature dimensions to $D/4$. All models process $224\times 224$ images during training and evaluation.
For pixel-level embedding, we use a stride-2 convolution, reducing feature map size by half. The positional feature map $\mathbf{F}_{pos}$ and superpixel index map $\mathcal{I}$ are down-scaled accordingly. A preliminary study showed that using stride 1 yielded negligible gains at a much higher computational cost, so we opted for stride 2.
Among the baseline methods, SPiT [1] is the most closely related to ours, as it also employs superpixels in tokenization. Meanwhile, Quadformer [49] and MSViT [19] do not utilize superpixels but improve the tokenization process through mixed-scale tokenization. Following the settings of each method, we conduct experiments in two different settings. For DeiT [55] and SPiT [1], we train and evaluate the models using the standard ImageNet-1K (IN-1K) [12] classification setting [55, 20, 39, 63, 68], where all models are trained from scratch. In contrast, for Quadformer [49] and MSViT [19], we follow their experimental setups. Specifically, MSViT initializes the transformer backbone with a ViT model pre-trained on ImageNet-21K (IN-21K) [48], while Quadformer uses a ViT model that is first pre-trained on IN-21K and then further fine-tuned on IN-1K. In our experiments, we follow the setting of Quadformer.
For the superpixel algorithm, we employ SLIC [2], but in practice, we use its faster implementation, FastSLIC [4] due to its computational efficiency. We configure SLIC [2] with its default setting, a compactness of 10, a maximum of 10 iterations, and 196 superpixels. Note that the value 196 is simply the maximum number of superpixels, and the actual number of superpixels varies across images. For example, on the ImageNet-1K [12] validation set, the average number of superpixels is approximately 191.9 ± 6.0.
5.2 Comparison to Baselines
Table 1 compares the classification performance of SuiT with baselines in two different settings. For each model size, the table provides a detailed comparison across key aspects: number of total parameters of the model (# Params.), computational cost measured in GMACs (GMACs), and the classification accuracy (Top-1).
The results in Table 1 demonstrate that SuiT consistently outperforms the baselines in both experimental settings. Specifically, when trained from scratch with randomly initialized weights, SuiT achieves performance gains across all model sizes compared to DeiT [55]: Tiny (+3.5%p), Small (+1.1%p), and Base (+0.5%p). Additionally, SuiT surpasses SPiT [1], the most closely related superpixel tokenization baseline, by substantial margins: Small (+5.9%p) and Base (+1.9%p).
In the setting where weights are initialized with IN-21K [48], SuiT consistently outperforms Quadformer [49] and MSViT [19] across all model scales. Notably, in the Tiny model, SuiT with random initialization achieves superior performance compared to MSViT [19] initialized with IN-21K [48] pre-trained weights. Furthermore, in the Small model, our approach trained with random initialization even surpasses Quadformer [49] initialized with IN-21K [48] + IN-1K [12] pre-trained weights.
5.3 Adaptive Inference
Figure 3 highlights the adaptive inference capability of SuiT. This figure illustrates the performance variation with respect to the number of tokens during inference, where SPiT [1] is plotted based on the average token count, while SuiT is plotted using the maximum token count. Unlike DeiT [55], which relies on a fixed number of tokens, SuiT can dynamically adjust the number of superpixel tokens during inference, enabling more efficient computation while maintaining strong performance. Notably, SuiT outperforms DeiT [55] even with fewer tokens, requiring only 100 tokens compared to DeiT’s 196 to achieve the same performance in the tiny model setting. This highlights its ability to extract meaningful visual representations with higher efficiency.
Additionally, SuiT can further increase the number of superpixels at inference time, effectively scaling accuracy as needed. Notably, despite being trained with 196 superpixels, increasing to 256 at inference improves accuracy across different model scales: from 75.7% to 75.9% in the Tiny model, 80.9% to 81.2% in the Small model, and 82.3% to 82.5% in the Base model. This contrasts with SPiT [1], which, despite its ability to adjust token counts, fails to deliver competitive performance across all token settings. In specific, SPiT-B [1] achieves only 64.0% accuracy with an average of 46 tokens, while increasing the token count to an average of 774 marginally improves performance to 79.2%. Note that SPiT [1] does not allow direct control over the number of tokens but instead relies on a “level” hyperparameter, limiting fine-grained adjustments and restricting evaluations to extreme settings (46 and 774 tokens).
These findings underscore the strength of SuiT’s adaptive inference, allowing it to flexibly balance computational cost and accuracy, making it a more scalable and efficient alternative to both DeiT [55] and SPiT [1].
5.4 Transfer Learning
We verify the generalization ability of SuiT when transferred to various downstream tasks. We follow the standard fine-tuning protocol and default settings of DeiT [55] for fair comparison, by fine-tuning the ImageNet-1K pre-trained SuiT to each downstream task specifically. We test our models on iNaturalist [57], Flowers102 [44], and StanfordCars [10]. Further details can be found in the Appendix A2.2.
@@ 5.4 Transfer Learning
Table 2: Results of transfer learning different datasets in terms of top-1 accuracy results. SuiT outperforms ViT in most downstream tasks in both supervised and self-supervised settings. : The reported Flowers102 performance could not be reproduced in the exact setting, so we compare with our reproduced results.
@@
Table 2 presents a comparison on the classification accuracy between DeiT and SuiT when fine-tuned on various downstream tasks. Across various tasks and model scales, SuiT consistently achieves higher accuracy than DeiT. Notably, iNaturalist2018 [25], a dataset of sufficient scale, SuiT-Small outperforms DeiT-Small by 0.8%p, highlighting its superior ability to transfer learned features effectively especially in data-rich scenarios.
5.5 Zero-shot Salient Object Segmentation
We conduct zero-shot salient object segmentation to demonstrate the advantages of superpixel tokens in dense prediction tasks. We adopt DINO [9] models for this experiment. We evaluate the performance on ECSSD [65], DUTS [61] and DUT-OMRON [66] using the TokenCut framework [64].
Superpixel tokenization significantly boosts DINO-ViT’s performance across most metrics, even without post-processing, as shown in Table 3. Unlike DINO-ViT, SuiT maintains minimal performance differences between processed and unprocessed results. Compared to SPiT [1], SuiT achieves competitive performance, though SPiT was trained on higher-resolution images.
Moreover, as Figure 4 shows, DINO-SuiT detects all salient objects, even in multi-object scenarios, whereas DINO-ViT struggles with rectangular masks and incomplete detections. The visualization of self-attention maps of DINO-SuiT can be found in the Appendix A2.3.
@@ 5.5 Zero-shot Salient Object Segmentation
5.6 Semantic Integrity of Superpixel Tokens
To assess the semantic preservation of our superpixel tokens, we perform K-means clustering [29] in the token embedding space and compare the results with those of patch tokens.
As shown in Figure 5, our superpixel tokens cluster based on semantic similarity, whereas DeiT’s patch tokens lack consistency. For example, in the first row, tokens from duckling regions are grouped together, while DeiT’s patch tokens fail to capture clear semantic meaning. This suggests that superpixel tokens better preserve semantic integrity.
SPiT, which also utilizes superpixels in tokenization, exhibits similar behavior to SuiT. However, SPiT clusters the shadows together with the ducklings in the first row, whereas those of SuiT selectively groups only the ducklings. This further supports that our tokenization method preserves stronger semantic coherence. Additional results can be found in the Appendix A3.
@@ 5.6 Semantic Integrity of Superpixel Tokens
Local feats. PE Freqs. PE Injection Top-1
random init. concat+proj 75.7
random init. concat+proj 66.5
No PE 71.5
pre-defined concat+proj 73.0
random init. concat 73.8
random init. addition 75.2
Avg Std Softmax Max Top-1
75.7
73.2
75.5
Fail
75.5
Compactness Iters Algorithm Top-1
10 10 Fast-SLIC 75.7
1 10 Fast-SLIC 74.4
100 10 Fast-SLIC 75.5
Square N/A N/A 75.1
10 1 Fast-SLIC 75.4
10 20 Fast-SLIC 75.6
0.001 N/A Watershed 73.9
Table 6: Ablation on the superpixel algorithm.
5.7 Ablation Studies
Pixel-level Embeddings
The following ablation studies on our proposed components are conducted using SuiT-Tiny. Table 4 highlights the importance of components in obtaining pixel-level embeddings. First, using raw 3-channel RGB features (row 2) leads to a significant performance drop, confirming the need for higher-dimensional local features. Second, removing positional information (row 3) reduces performance but retains partial class prediction through local cues. Experiments also reveal that pre-defined positional frequencies (row 4) under-perform compared to learned frequencies, emphasizing the need for vision-specific customization. Lastly, our results show that the combination of concatenation and projection more effectively integrate local and positional features than simple concatenation (row 5) or addition (row 6) used in prior studies [58, 14, 55].
Aggregation
Table 5 presents the analysis of our aggregation method, which employs cardinality-independent operations. Removing max pooling (row 2) results in a notable performance drop, and removing average pooling (row 3) causes a slight decrease. We also tested standard deviation as an alternative to max pooling, but it proved numerically unstable and unsuitable for training (row 4). Softmax pooling (row 5), which computes a weighted sum of values within each superpixel, showed promising results but slightly underperformed compared to max pooling. This may be due to the softmax operation suppressing prominent features crucial for recognition, leading to less discriminative representations.
Superpixel Method
In Table 6, we present a study regarding the superpixel methodology. First, we examine the compactness parameter in SLIC, which controls the trade-off between spatial proximity and pixel content in superpixel formation (rows 2-4). Higher values produce more regular shapes but reduce semantic relevance. We compare compactness values of 1 (row 2) and 100 (row 3), along with square-shaped tokens (row 4) as an upper-bound of high compactness. Keeping other components unchanged, we find that increasing compactness degrades performance, supporting the importance of semantic integrity. However, an extremely low value (e.g., 1) also harms performance, likely due to insufficient spatial coherence. Our framework relies on SLIC, which iteratively clusters pixels. We try different number of iterations in clustering as shown in Table 6, but the number of iterations had little impact on performance (rows 5 and 6). We also test Watershed [43] (row 7), which results in a slight performance drop but remains robust.
6 Limitation and Future Work
Our method, in its current form, is limited to the vanilla ViT architecture[14]. Yet, since the ViT architecture is the most prevalent choice for vision architecture in many applications, e.g., CLIP [47] and LLaVA [37] in multi-modal learning, we believe that building an improvement for the ViT architecture is valuable. Applying superpixel tokenizations to other ViT variants, e.g., Swin Transformer [39] remains a topic for future study.
7 Conclusion
In this work, we proposed a novel superpixel-based tokenization method for ViTs as an attempt to improve the visual tokenization process. By leveraging superpixels, the proposed tokenization obtains tokens preserving semantic integrity, overcoming the issue of mixed semantics in grid-based patches. To incorporate superpixels that are irregular in shape, size, and location, we proposed a two-stage pipeline consisting of pixel-level embedding and superpixel-aware aggregation. Extensive experiments demonstrated the effectiveness and intriguing properties of our method. We hope our promising results inspire further exploration of token design in ViTs, fostering advancements in both model performance and interpretability.
Supplementary Material
Figure A1: Class-identifiability analysis results of DeiT and SuiT. We show the proportion of class-identifiable [CLS] tokens and image tokens for SuiT and DeiT across all layers.
@@ 7 Conclusion
Table A1: Experimental results of SuiT and DeiT on ImageNet-A and ImageNet-O. We report accuracy on ImageNet-A and AUPR on ImageNet-O, following  [24].
A1 Additional Experiments
A1.1 Class Identifiability Analysis
[59] examined class-specific encoding in supervised ViTs by projecting intermediate representations through the trained classifier head into a class-specific embedding space, similar to the logit lens method in NLP [45, 6]. Building on this, we evaluate class-specific encoding across SuiT’s layers by projecting token representations through the classifier head and measuring the proportion of tokens correctly predicting the target class on IN-S-919 [18].
Figure A1 shows that SuiT’s classification token shows higher class identifiability in the earlier layers compared to DeiT. This can be attributed to the semantic preserving nature of superpixel tokenization, which aids in aggregating class-specific encoding in early layers. In the deeper layers, DeiT’s patch tokens exhibit class identifiability levels slightly lower than its classification token, with a marginal gap. This observation aligns with findings from previous studies [59, 60], which report that patch tokens in supervised ViTs tend to encode increasingly global information as layers deepen. In contrast, SuiT’s patch tokens maintain a lower proportion of class-identifiable features in the later layers, suggesting a reduced emphasis on global feature encoding. This divergence alludes that SuiT and DeiT adopt distinctive strategies to aggregate class-specific features at different network depths.
A1.2 Enhanced Robustness
@@
In Table A1, SuiT consistently outperforms DeiT on both datasets across model variants. These findings indicate that our superpixel tokenization improves model robustness in challenging scenarios by mitigating the model’s reliance on features that are easily learned but lack generalizability.
A2 Implementation Details
A2.1 ImageNet-1K classification
A2.1.1 Training from scratch
The settings for experiments from scratch is presented in Table A3. We mostly follow the settings of DeiT [55] with minimal changes. The main difference to the DeiT setting is as follows: 1) to align with our main concept of using superpixels, we disable MixUp [70]. 2) In random erasing [73], the randomly erased region is filled with per-channel random color instead of per-pixel random color. 3) We use gradient clipping, which was known to be helpful in ViT [14], but later removed in DeiT [55]. Other than that, we slightly tweak hyper-parameters of learning rate, weight decay and stochastic depth. For models of all three scales, Tiny, Small and Base, we use a base learning rate of $48e-$ 4, which is scaled proportionally to the batch size, by ×$\times\frac{batchsize}{512}$. We use different weight decay and stochastic depth values for each scale. Most notably, our model uses a larger weight decay and drop rate for stochastic depth, especially in Base scale. This is because we found that our model tends to overfit quickly, compared to the baseline DeiT. We thus opt to use a stronger regularization term. Similar observations were made in Transfer Learning, which is also discussed in the Section  A2.2.
A2.1.2 Fine-tuning from Pre-trained weights
@@
Figure A2: Self-attention map comparison between DINO-ViT and DINO-SuiT. All maps are visualized by averaging the [CLS] token’s self-attention from the final layer across all heads.
A2.2 Transfer Learning
@@
A2.3 Self-Supervised Learning
For DINO [9] pre-training, we follow the exact training hyper-parameters used to pre-train ViTs, without any tuning, extra training techniques or augmentations. We train DINO-SuiT for 300 epochs, and transfer the pre-trained model to downstream tasks following the settings mentioned in Section A2.2. Additional self-attention maps of DINO-SuiT can be found in Figure A2 and A4.
Figure A3: Visulization of superpixel tokenization using various parameters.
A2.4 Zero-shot Salient Object Segmentation
@@
A3 Semantic Integrity of Superpixel Tokens
@@
A4 Visualization of Superpixel Tokenization
We visualize how an image can be tokenized using various parameters in superpixel generation process at Figure A3. We show how the tokens change when setting different number of superpixels to be generated (compactness set to 10) and when setting different number of compactness (number of superpixels set to 196).